OpenStudy (anonymous):

Which expression is equal to (5-2i)-(1+3i) A)4+3i B)4+i C)6-5i D)4-5i i think it's A

OpenStudy (anonymous):

So you'd simplify this exactly like you would multiply out brackets.

OpenStudy (anonymous):

I don't even know how to multiply out brackets.

OpenStudy (lynfran):

u dont have to multiply out the brackets

OpenStudy (lynfran):

just distribute the negative sign to the second bracket

OpenStudy (lynfran):

so we have 5-2i-1-3i

OpenStudy (lynfran):

now group like term and simplify

OpenStudy (lynfran):

and no its not A

OpenStudy (anonymous):

I only thought it was A because if you do 5-2=4 and 1+3=4

OpenStudy (lynfran):

but 5-2i arent like term... they can not =4

OpenStudy (lynfran):

5-2i-1-3i = 5-1 are like term and -2i-3i are like terms

OpenStudy (anonymous):

so it would be 4 and -5

OpenStudy (lynfran):

4-5i

OpenStudy (anonymous):

Thank you could help me with one more?

OpenStudy (lynfran):

ok

OpenStudy (anonymous):

Which expression is equal to -2i(4-i) A)8-2i B)-2-8i C)-8-2i D)2-8i

OpenStudy (lynfran):

distribute the -2i to the bracket

OpenStudy (anonymous):

you do 2x4=8i so it would be 2-8i??

OpenStudy (lynfran):

and none of those expressions would equal to that you shd get -2i*4+2i^2 -8i+2i^2

OpenStudy (anonymous):

I'm a bet confuse on what should I do??

OpenStudy (lynfran):

-2i multiple to 4 = -8i and -2i multiply to -i =2i^2

OpenStudy (lynfran):

r u sure u wrote the expression correctly??